Andy, Eli, and I went out on the night of the 4th. Temperature was around 65F, focus set to 2865, and FWHM varied between 4.25 and 7 pixels. We looked at the open cluster IC 4665 in 4 sections since it covers a lager area of the sky than the telescope can capture in one shot. We also took lots of extra flats and bias images that Eli will use for linearizing the telescope and his other research.
We had some trouble with DFM TCS and Maxim DL crashing during the night and the computer froze at the end of the night while we were taking the darks so we had to restart the system. We also could not get the telescope to return to it’s starting position at the end of the night. We tried restarting the computer and DFM TCS but the normal procedures wouldn’t work so we put the cover on and had to leave it as it was.